# Nonlinear Fore(Back)casting and Innovation Filtering for Causal-Noncausal VAR Models

### v3

@@ 3.2 Out-of-Sample Predictive Density
where $l_{2}(z_{2})$ is the stationary pdf of $Z_{2,t}=A^{2}\left(\begin{array}[]{c}Y_{t}\\ \tilde{Y}_{t-1}\end{array}\right)$, if $n_{2}\geq 1$. In the pure noncausal process, $n_{2}=0$, we have $l(y|\underline{Y}_{T})=g(y-\Phi_{1}Y_{T}-\cdots-\Phi_{p}Y_{T-p+1}).$
@@
This predictive density is constrained by the VAR representation (3.1). It is a semi-parametric function of parameters $\Phi_{1},...,\Phi_{p}$, determining $A^{2}$ and $J_{2}$, and of functional parameters $g,l_{2}$. Except if the distribution is multivariate Gaussian, the predictive density has a complicated form, with several modes and a conditional mean which is not linear in $\underline{Y}_{t-1}$, in general. From equation (3.12) and the deterministic relation between $Y_{t}$ and $Z_{t}$, it follows that:
Corollary 1: Markov Property
The mixed VAR(p) process $(Y_{t})$ [resp. the state process $(Z_{t})$] is a Markov processes of order $p$ [resp. of order 1] in calendar time for $n_{2}\geq 1$. They are both Markov in reverse time too.
This corollary extends to linear mixed causal-noncausal processes of any autoregressive order, the result of Cambanis, Fakhre-Zakeri (1994), who show that a linear pure noncausal autoregressive process of order 1 is a causal Markov process of order 1.
@@ 3.3 Backward Predictive Density
Since the mixed VAR process of order $p$ is Markov of order $p$ both in calendar and reverse time, we can derive the closed-form expression of the backward predictive density in reverse time for backcasting.
Corollary 2: Backcasting
In a mixed VAR(1) model the backward predictive density of $Y_{T-1}$ given $Y_{T}$ is:
@@
Proof: See Appendix A.1.
3.4 Discussion
@@
By considering jointly Proposition 1 and Corollary 2, we can see that the mixed VAR models have a nonlinear dynamic structure that allows for extending to nonlinear framework the standard Kalman filter for linear Gaussian processes.
@@
We observe that there exists a multiplicity of real Jordan representations and of matrices $A$ built from extended real eigenspaces (in the presence of complex conjugate eigenvalues). However, $\det J_{2}=\prod_{j=1}^{n_{2}}\lambda_{j}$, where $|\lambda_{j}|>1,\,j=1,...,n_{2}$, is independent of the real Jordan representation. Similarly, the noncausal component $Z_{2}$ is defined up to a linear invertible transformation. Since the Jacobian is the same for the numerator and denominator of the ratio $\frac{l_{2}\left[A^{2}\left(\begin{array}[]{c}y\\ \tilde{Y}_{T}\end{array}\right)\right]}{l_{2}\left[A^{2}\left(\begin{array}[]{c}Y_{T}\\ \tilde{Y}_{T-1}\end{array}\right)\right]}$, it has no effect on the ratio. Thus, the expression of $l(y|\underline{Y}_{T})$ does not depend on the selected real Jordan representation, that is on the selected state-space representation.
A mixed VAR process has nonlinear causal dynamics, which is captured by the predictive density. For this nonlinear and non-Gaussian process, the predictive density provides the oos prediction intervals at various horizons, and replaces the linear pointwise predictions and prediction intervals of causal VARs.
@@
The closed-form expression of the predictive density is available at horizon 1. It can be written at horizon $h$ as a multivariate integral over $h$ future values of the process. For given $\Phi_{1},..,\Phi_{p}$ and $g$, the predictive density at horizon $h$ can be approximated as follows: The closed-form predictive density at horizon 1 allows us to perform drawings of future values of the process, by using the Sampling Importance Resampling (SIR) method [Gelfand, Smith (1992), Tanner (1993)] (see Section 6.1.4 for an illustration). To approximate the predictive density at horizon $h>1$, for given $\Phi_{1},\Phi_{2},...$ and $g$, we need S independent future paths of the process. Each path $s=1,...,S$ is obtained by forecasting sequentially $Y_{T+1}^{s}|Y_{T}$, followed by $Y_{T+2}^{s}|Y_{T+1}^{s},...,Y_{T+h}^{s}|Y_{T+h-1}^{s}$, from the predictive densities. This provides a drawing $Y_{T+1}^{s},...,Y_{T+h}^{s}$ of a future path $s$. By replicating independently for $s=1,...,S$, we get $Y_{T+h}^{s},\;s=1,...,S$ and can use their sample distribution as an estimator of the predictive distribution at horizon $h$.
4 Nonlinear Causal Innovations
In macroeconomics and finance, the use of the traditional causal VAR model commonly includes the IRF analysis, which is an important tool for the economists and financial policy makers. These tools are used for the so-called causal analysis. This terminology differs from the ”causal-noncausal” terminology introduced by Rosenblatt in the context of time series, although they are related. More precisely, the ”causal analysis” implies a) the possibility to make inference on the future from the past, i.e. to make forecasts and also b) the counterfactual analysis of the effects of current transitory shocks on the future. The standard linear causal IRF analysis cannot be applied to the mixed VAR(p) processes, because neither errors $\varepsilon_{t}$ in model (2.1), nor state-specific linear innovations $\eta_{t}$ in (3.8) are causal innovations independent of the lagged values of $Y_{t}$  [^6] Some components of $\varepsilon_{t}$ depend on the past values of $Y_{t}$. Then, it is difficult to interpret the shock on $\varepsilon_{t}$ that implicitly changes a past that has already been realized [see Davis, Song (2020), Figures 1 and 7 for this practice].. Moreover the presence of noncausal roots in the mixed VAR model implies nonlinear causal dynamics, characterized by bubbles, for example. The nonlinearity has to be accounted for in the definition of causal innovations. In addition, these nonlinear causal innovations have to satisfy both the serial and cross-sectional independence conditions [Gourieroux, Jasiak (2005), Gourieroux, Monfort, Renne (2017)]. As mentioned earlier, the errors $\varepsilon_{t}$ of mixed VAR models cannot be interpreted as innovations and then used to define shocks and IRFs. We discuss below the filtering and identification of nonlinear causal innovations in the framework of mixed VAR models.
4.1 Definition of Nonlinear Innovations
Definition 1: A nonlinear causal innovation of process $(Y_{t})$ is a process $(v_{t})$ of dimension $m$ such that: i) the vectors $v_{t}$ are serially i.i.d.; ii) the strictly stationary process $(Y_{t})$ can be written in a nonlinear autoregressive form:
$Y_{t}=a(\underline{Y_{t-1}},v_{t}).$ (4.1)
iii) the future values $\bar{v}_{t}$ of the innovation process are independent of the lagged values $\underline{Y_{t-1}}$ of the observed process $Y_{t}$.
It is easy to see that conditions i) and ii) are equivalent to the Markov of order $p$ property of $(Y_{t})$ with a continuous distribution, and can be applied to the mixed model (3.1) by Corollary 1 [see Appendix A.3 a)]. If function $a$ is invertible with respect to $v$, then $v_{t}$ is a nonlinear function of $Y_{t}$ given its past and condition iii) is satisfied. Under conditions i) and ii) the autoregressive equation (4.1) can be applied recursively.
The nonlinear autoregressive model (4.1) is a nonlinear causal representation of the mixed VAR. The sequence of nonlinear causal innovations provide the basis of nonlinear impulse response functions [Potter (2000), Gourieroux, Jasiak (2005)].
Let us consider a transitory shock $\delta$ at date $T$ on $v_{T}$, and then apply recursively equation (4.1) to get the impulse response function for the future of the process:
$Y_{T}^{\delta}=a(Y_{T-1}^{\delta},v_{T}+\delta)$, $Y_{T+1}^{\delta}=a(Y_{T}^{\delta},Y_{T-1},...,v_{T+1})$, and so on. The conditions i) and iii) in Definition 1 are crucial for the interpretation of shocks in terms of ”causal analysis”. Condition iii) means that $v_{T}$ can be shocked at any time T without an effect on the realized past $\underline{Y}_{T-1}$. Condition i) means that this shock has no effect on the future values $\bar{v}_{T+1}$.
The nonlinear IRF defined above corresponds to a multivariate shock $\delta$. The literature on ”causal analysis” is also interested in univariate specific shocks. It is possible to define shocks on the component $v_{1,t}$ of $v_{t}$, if additionally the component $v_{1,t}$ and ($v_{2,t},...,v_{n,t}$) are cross-sectionally independent.
4.2 Identification of Nonlinear Autoregression and Innovation
The nonlinear autoregression (4.1) depends on function $a$ and on the distribution of the causal nonlinear innovation $v_{t}$. In the mixed VAR(p) model, these both depend on $\Phi_{1},\Phi_{2},...$ and $g$, but do not satisfy a one-to-one relationship with them because the nonlinear function $a$ depends on much more arguments than the density $g$.
This new nonparametric identification issue is examined in Appendix A.3 in a functional framework by using the properties of harmonic functions. In particular, we get the following Proposition:
Proposition 2:
In a mixed VAR(1) process, the dimension of under-identification in the functional space of nonlinear autoregressive models is finite and equal to $2m$.
This functional multiplicity is very large. Appendix A.3 shows that it is possible to restrict the analysis to nonlinear autoregressions with Gaussian innovations, such that the components $v_{1,t},...,v_{n,t}$ are independent N(0,1). However, this additional restriction is not sufficient to solve the identification issue. Indeed, there exist nonlinear transformations of $v_{t}$, with respect to which the multivariate N(0, Id) distribution remains invariant. As shown in Appendix A.3, they correspond to ”local matrix rotations”, i.e. rotations depending on the level $v$.
Corollary 3: Functional multiplicity
There exists a functional multiplicity of innovation processes ($v_{t}$) that are cross-sectionally independent standard normal variables.
Since the observed process ($Y_{t}$) and the state process ($Z_{t}$) satisfy a one-to-one relationship, for structural interpretations discussed below it is often preferable to consider nonlinear causal innovations to the state variables. The state variables satisfy also a nonlinear autoregressive scheme:
$Z_{t}=\tilde{a}(\underline{Z}_{t-1},\tilde{v}_{t}),$ (4.2)
where the $\tilde{v}_{t}$ can be serially conditionally i.i.d. variables (possibly restricted to be standard Gaussian) .
Therefore, the identification of nonlinear autoregressive representation of Gaussian nonlinear innovations and of nonlinear IRFs is a major issue in the mixed VAR models.
Example 1: The SVAR Model
For the structural causal VAR (SVAR) models, a similar identification issue has been recently solved [see, Online Appendix B]. More precisely, let us assume a causal SVAR model written as: $Y_{t}=\Phi_{1}Y_{t-1}+...+\Phi_{p}Y_{t-p}+\varepsilon_{t},\;\;\varepsilon_{t}=Du_{t},$ where the error term $\varepsilon_{t}$ is a linear transform of independent sources $u_{t}$, with distributions $g_{1},..,g_{n}$, respectively. Then, if at most one source is Gaussian, it is possible to identify $\Phi_{1},...,\Phi_{p},D,g_{1},...,g_{n}$ [see, Comon (1994) for ICA, Gourieroux, Monfort, Renne (2017), (2020), Guay (2021), Bernoth et al. (2024) for the application to SVAR models]. Then the nonlinear autoregressive model is:
$Y_{t}=\Phi_{1}Y_{t-1}+...+\Phi_{p}Y_{t-p}+D\left(\begin{array}[]{c}G_{1}^{-1}[\Phi(v_{1,t})]\\ \vdots\\ G_{1}^{-1}[\Phi(v_{n,t})]\end{array}\right),$
where the components $v_{i,t}$ are independent standard normal variables, uniquely defined from $v_{i,t}=\Phi^{-1}[G_{i}(u_{i,t})]$, where $G_{i}$ is the c.d.f. of $g_{i}$ and $\Phi$ is the c.d.f. of the standard normal. Therefore the autoregressive function is the sum of a linear function of $\underline{Y}_{t-1}$ and a nonlinear function of $v_{t}$. In macroeconomic applications, shocks with economic interpretations, such as the monetary shock, fiscal shock or productivity shock are usually applied to specific linear combinations $c^{\prime}Y_{t}$ of the observed variables, which play the role equivalent to our state variables $z_{t}=c^{\prime}Y_{t}$.
There exists a variety of methods to circumvent the functional identification issue.
i) In the applied literature, identifying restrictions of function $a$ are often introduced by means of parametric assumptions on the joint distribution of errors $\varepsilon_{t}$. For example,one could choose this distribution among non-Gaussian elliptical parametric families based on t-Student distributions. However, such a partial identifying assumption is difficult to interpret due to the lack of structural interpretation of errors ($\varepsilon_{t}$).
ii) In special cases, a specific structural shock can appear. For example, the applications of mixed models indicate that the noncausal order is often equal to 1, i.e. $n_{2}=1$ [see e.g. Gourieroux, Jasiak (2017), Hecq, Lieb, Telg (2016), and Section 6.3 of this paper]. This is likely due to the type of nonlinear dynamics generated by noncausal roots capturing the speculative bubbles and local trends. For example, in macroeconomic models, the noncausal component can be related to speculative bubbles in oil prices impacting jointly the price index, GDP and other macroeconomic variables. It is also common in financial applications based on nonlinear models with one factor capturing the nonlinear dynamics and interpreted as a systemic component. This explains the recent interest in common bubble [Gourieroux, Zakoian (2017), Cubbada et al. (2023), Hall, Jasiak (2024)].
In the case $n_{2}=1$, with the interpretation of $Z_{2}$ as the systemic factor, it is insightful to examine the consequences of a shock to the latent noncausal component $(Z_{2})$, which determines the nonlinear dynamics of the model rather than to the observed variables. This can be done by applying nonlinear shock ordering to the pure causal and noncausal components $(Z_{t})$ and selecting $Z_{2}$ as the first variable to be shocked. To simplify the notation, we write $v_{t}$ defined in equation (4.2) without the tilde. Then the nonlinear Gaussian causal innovation of $Z_{2}$ is uniquely defined as:
$v_{2,t}(Z_{t})=\Phi^{-1}[F_{2}(Z_{2,t}|\underline{Y}_{t-1})]=\Phi^{-1}[F_{2}(Z_{2,t}|Z_{t-1})],$ (4.3)
where $F_{2}$ is the conditional cumulative distribution function (c.d.f) of $Z_{2,t}$. From equation (a.5) given in Appendix A.1, it follows that the conditional density of $Z_{2,t}$ given $Z_{t-1}$ has a closed form given by:
where $g_{\eta_{2}}$ is the marginal density of $\eta_{2}$. The c.d.f. $F_{2}$ is the integral of $l(z_{2,t}|z_{t-1})$ over the set of admissible values of $z_{2,t}$.
Next, we can append it by searching for a possible Gaussian ($v_{2,t},...v_{n,t}$) associated with the causal state variables. This can be done recursively [see Appendix A.3]. For illustration, let us consider the case $n_{1}=1$ and define:
$v_{1,t}(Z)=\Phi^{-1}[F_{1|2}(Z_{1,t}|Z_{2,t},Z_{t-1})],$ (4.5)
where $F_{1|2}$ is the conditional cumulative distribution function of $Z_{1,t}$ given $Z_{2,t},Z_{t-1}$. From equation (a.5) in Appendix A.1, we get the closed-form expression of the conditional density:
When the variable $Z_{2}$ defines the selected ”ordering” it is easy to check that $v_{1,t}(Z)$ differs in general from the causal innovation $\eta_{1,t}$ to the latent causal component $Z_{1,t}$. By applying the ordering, we see from (4.3)-(4.5) that the autoregressive model can be written under a recursive form. Moreover, we deduce from (4.4):
Corollary 4:
The noncausal state variable $(Z_{2,t})$ is a Markov process of order 1 with respect to the filtrations associated with $(Y_{t})$, $(Z_{t})$ and $(Z_{2,t})$.
By construction, the structural noncausal shock $v_{2,t}(Z_{t})$ is defined in a unique way (since $n_{2}=1$) and is a Gaussian white noise  [^7] See Gourieroux, Jasiak (2005) for the uniqueness of nonlinear Gaussian innovation in a univariate model.. It is also independent of the multivariate shock $v_{1,t}(Z_{t})$. Then, we can construct nonlinear IRFs for $Y$ based on the effect of a change $\delta_{2}$ in $v_{2,t}$, with $v_{1,t}$ held constant. It is called the Common Bubble Shock (CBS) henceforth. In the general case $n_{2}=1,n_{1}\geq 1$, the associated multivariate impulse responses of $Z_{2,t}$ are identifiable, i.e. independent of other components ($v_{1,t},v_{3,t},...v_{n,t}$).
@@ 5 Statistical Inference
The parameters of a mixed VAR model need to be estimated before the forecasts and nonlinear causal innovations are computed. The first part of this Section reviews the estimation methods that exist in the literature and describes the new prediction/filtering algorithm for the estimated nonlinear causal innovations. Next, we introduce a post-estimation inference method for assessing the prediction interval uncertainty due to the preliminary estimation step.
@@ 5.1 Estimation and Filtering
The mixed VAR model can be estimated by the maximum likelihood method based on an assumed parametric distribution of $\varepsilon_{t}$ [see Breidt et al. (1991), Lanne, Saikkonen (2010), (2013), Davis, Song (2020), Bec et al. (2020)]. This approach yields consistent estimators provided that the parametric distributional assumption is well specified.
Alternatively, the mixed VAR model can be consistently estimated without any parametric assumptions on the distribution of the errors  [^8] Except for the non-Gaussianity assumption. by using the (Generalized) Covariance (GCov) estimator [Gourieroux, Jasiak (2022)]. The GCov estimator is consistent, asymptotically normally distributed and semi-parametrically efficient. As an alternative, minimum distance estimators based on the cumulant spectral density of order 3 and 4 have been proposed in Velasco, Lobato (2019) and Velasco (2022) 999 See also Starck (2023) for a two-step continuum method of moments. The method of moment estimators provide consistent estimators only in applications to the VAR models satisfying the additional cross-sectional independence condition [see, Guay (2021)]..
The prediction and filtering methods introduced in Sections 3,4 for given $\Phi_{1},...,\Phi_{p}$ and $g$ can be applied in a parametric or semi-parametric framework by replacing the unknown parameters by their consistent estimates. In the semi-parametric framework, it can be applied along the following lines:
@@
step 2. Use the $\hat{\Phi}_{i},i=1,...,p,$ estimates to compute the roots of the lag-polynomial and more generally an estimated Jordan representation: $\hat{A},\hat{J}_{1},\hat{J}_{2}$.
@@
- the density $l_{2}$ of $Z_{2,t}$ can be estimated from $\hat{Z}_{2,t},\;t=1,...,T$;
step 6. The predictive density can be estimated from the formula of Proposition 2 by replacing $l_{2}$ by $\hat{l}_{2}$, $A^{2}$ by $\hat{A}^{2}$, and also $J_{2}$ by $\hat{J}_{2}$, $g$ by $\hat{g}$, and $\Phi_{1},...,\Phi_{p}$ by $\hat{\Phi}_{1},...,\hat{\Phi}_{p}$. The mode (median) of the predictive density provides the point forecasts and the quantiles of the predictive density can be used to obtain estimated prediction intervals at horizon 1.
When there is a single non-causal state variable this can be completed by:
step 7. The approximations of nonlinear causal innovations $\hat{v}_{2,t}(Z)$ can be computed from the estimated distributions as $\hat{v}_{2,t}(Z)=\Phi^{-1}(\hat{F}_{2,T}(\hat{Z}_{2,t}|\hat{Z}_{t-1}))$ by applying the formula of predictive density (4.3) with $l_{2}$ replaced by $\hat{l}_{2}$ and $g_{\eta_{2}}$ replaced by $\hat{g}_{\eta_{2}}$, i.e. the empirical density of $\hat{\eta}_{2,t},\,t=1,...,T$.
step 8. Next, they can be appended by the approximated nonlinear causal innovations independent of $\hat{v}_{2,t}(Z)$ such as
$\hat{v}_{1,t}=\Phi^{-1}(\hat{F}_{1|2,T}(\hat{Z}_{1,t}|\hat{Z}_{2,t},\hat{Z}_{t-1})),\;\mbox{for}\;n_{1}=1$.
5.2 Uncertainty of the Estimated Prediction Set
The estimated model parameters and residuals $\hat{\varepsilon}_{t}$ can be used to build oos forecast intervals conditional on given values of the last observations in the sample, called the conditional prediction interval. The estimation errors associated with the scalar and functional parameters have an effect on the uncertainty of the conditional prediction interval. We provide below a detailed framework in which various notions of prediction sets can be distinguished.
i) True Prediction Interval
For ease of exposition, let us consider the VAR(1) model, forecast horizon $h=1$ and future value of the first component series $Y_{1,T+1}$ to be forecast at date $T$ out of sample (oos) given $Y_{T}=(y_{1,T},y_{2,T})^{\prime}\equiv y$, where $(Y_{2,t})$ contains the remaining components of the series. Then, the true prediction interval at level $1-\alpha_{1}$ for $Y_{1,T+1}$ is:

### v4

@@ 3.1 Forward Predictive Density
where $l_{2}(z_{2})$ is the stationary pdf of $Z_{2,t}=A^{2}\left(\begin{array}[]{c}Y_{t}\\ \tilde{Y}_{t-1}\end{array}\right)$, if $n_{2}\geq 1$. In the pure noncausal process, $n_{2}=0$, we have: $l(y|\underline{Y}_{T})=g(y-\Phi_{1}Y_{T}-\cdots-\Phi_{p}Y_{T-p+1}).$
@@
This predictive density is defined for the VAR representation (2.1). It is a semi-parametric function of the parameters $\Phi_{1},...,\Phi_{p}$, determining $A^{2}$ and $J_{2}$, and of the functional parameters $g,l_{2}$. This predictive density is complicated, except if it is multivariate Gaussian. In general, the conditional mean is not linear in $\underline{Y}_{T}$.
From equation (3.1), the deterministic relation (2.4) between $Y_{t}$ and $Z_{t}$, and the symmetry of the calendar and reverse time scales, it follows that:
Corollary 1: Markov Property The mixed VAR(p) process $(Y_{t})$ of dimension $m$ [resp. the state process $(Z_{t})$ of dimension $n=mp$] is a Markov processes of order $p$ [resp. of order 1] in calendar time for $n_{2}\geq 1$. The processes $(Y_{t})$ and $(Z_{t})$ are both Markov of orders $p$ and 1, respectively, in reverse time too.
This corollary extends to any autoregressive order $p$ the result of Cambanis and Fakhre-Zakeri (1994), who show that a linear pure noncausal autoregressive process of order 1 is a causal Markov process of order 1. It also extends the Proposition 3.1 in Freis and Zakoian (2019) derived in the univariate case.
The predictive density summarizes the nonlinear causal dynamics of a mixed VAR process. From the predictive densities, we infer the point forecasts and prediction intervals based on its quantiles. These lead us to the oos point predictions and prediction intervals at various horizons for this nonlinear and non-Gaussian process analogous to the linear pointwise predictions and prediction intervals of the causal VAR models.
@@ 3.2 Backward Predictive Density
Since the mixed VAR process of order $p$ is Markov of order $p$ both in calendar and reverse times, we can derive from Proposition 1 the closed-form expression of backward predictive density in reverse time for backcasting. For ease of exposition, we present it below for $p=1$.
Corollary 2: Backcasting Let us consider a mixed VAR(1) model. The backward predictive density of $Y_{T-1}$ given $Y_{T}$ is:
@@
Proof: See Online Appendix A.1.
@@
By considering jointly Proposition 1 and Corollary 2, we see that the mixed VAR models have a nonlinear dynamic structure extending to nonlinear dynamic framework the standard Kalman filter available for linear Gaussian processes.
@@
In Section 2.3, we mentioned that there exists a multiplicity of real Jordan representations of matrix $\Psi$. It implies a multiplicity of matrices $A$ built from its extended real eigenspaces (in the presence of complex conjugate eigenvalues). However, $\det J_{2}=\prod_{j=1}^{n_{2}}\lambda_{j}$, where $|\lambda_{j}|>1,\,j=1,...,n_{2}$, is independent of the real Jordan representation. Similarly, the noncausal component $Z_{2}$ is defined up to a linear invertible transformation. Since the Jacobian is the same for the numerator and denominator of the ratio $l_{2}\left[A^{2}\left(\begin{array}[]{c}y\\ \tilde{Y}_{T}\end{array}\right)\right]\;/\;l_{2}\left[A^{2}\left(\begin{array}[]{c}Y_{T}\\ \tilde{Y}_{T-1}\end{array}\right)\right]$, it has no effect on the ratio. Thus, the expression of $l(y|\underline{Y}_{T})$ does not depend on the selected real Jordan representation, that is on the selected state-space representation. The same remark applies to the backward predictive density.
3.3 Prediction at Horizon $h$
@@
The closed-form expressions of the backward and forward predictive densities at horizon 1 can be used sequentially to forecast or back-cast out-of-sample at any horizon $h>1$. Alternatively, the predictive densities at horizons $h>1$ can be obtained by using the Sampling Importance Resampling (SIR) method. Then, the forward predictive density at horizon $h$ can be written as a multivariate integral over $h$ future values of the process and approximated for given $\Phi_{1},..,\Phi_{p}$ and $g$ by drawing the future values of the process from the closed-form predictive density at horizon 1 replicated $S$ times by the SIR method [Gelfand and Smith (1992), Tanner (1993)] (see Section 6 for an illustration). Specifically, to approximate the predictive density at horizon $h>1$, for given $\Phi_{1},\Phi_{2},...$ and $g$, we need S independent future paths of the process. Each path $s=1,...,S$ is obtained by forecasting sequentially $Y_{T+1}^{s}|Y_{T}$, followed by $Y_{T+2}^{s}|Y_{T+1}^{s},...,Y_{T+h}^{s}|Y_{T+h-1}^{s}$, from the predictive densities. This becomes a drawing $Y_{T+1}^{s},...,Y_{T+h}^{s}$ of a future path $s$. By replicating it independently for $s=1,...,S$, we get $Y_{T+h}^{s},\;s=1,...,S$. Then, for a large $S$, we can use the sample distribution of $Y_{T+h}^{s},\;s=1,...,S$ as an estimator of the predictive distribution at horizon $h$.
@@ 4 Statistical Inference
The parameters of a mixed VAR model need to be estimated before the forecasts are computed. Below, we review the existing estimation methods in the time domain and describe the predictive algorithm providing the estimated predictive densities in a semi-parametric setup. As pointed out in Section 3, from the quantiles of the estimated predictive density we obtain the point and interval forecasts, which can be impacted by the preliminary estimation step involving estimators converging at different rates. Therefore, in the second part of this section we study the forecast interval uncertainty, using the theory of random sets [see Molchanov and Molinari (2018)].
@@ 4.1 Estimation and Filtering
The mixed VAR model can be estimated by the maximum likelihood method based on an assumed parametric distribution of $\varepsilon_{t}$ [see Breidt et al. (1991), Lanne and Saikkonen (2013), Davis and Song (2020)]. This approach yields consistent estimators provided that the parametric distributional assumption is correct (and non-Gaussian for identification).
Alternatively, the mixed VAR model can be consistently estimated without any parametric assumptions on the error distribution  [^4] Except for the non-Gaussianity assumption and the uniform tail parameter. by using the semi-parametric (Generalized) Covariance (GCov) estimator [Gourieroux and Jasiak (2023)]. The GCov estimator is consistent, asymptotically normally distributed and semi-parametrically efficient. As an alternative in the frequency domain, minimum distance estimators based on the cumulant spectral density of order 3 and 4 have been proposed in Velasco and Lobato (2018) and Velasco (2022).
The prediction methods introduced in Section 3 for given $\Phi_{1},...,\Phi_{p}$ and $g$ can be applied in a parametric or semi-parametric framework by replacing these unknown parameters by their consistent estimates. In the semi-parametric framework, this can be done along the following lines:
@@
step 2. Use the $\hat{\Phi}_{i},i=1,...,p,$ estimates to compute the roots of the lag-polynomial and more generally an estimated real Jordan representation: $\hat{A},\hat{J}_{1},\hat{J}_{2}$.
@@
- the density $l_{2}$ of $Z_{2,t}$ can be estimated from $\hat{Z}_{2,t},\;t=1,...,T$ (see Online Appendix B);
step 6. The predictive density can be estimated from the formula (3.1) by replacing $l_{2}$ by $\hat{l}_{2}$, $A^{2}$ by $\hat{A}^{2}$, and also $J_{2}$ by $\hat{J}_{2}$, $g$ by $\hat{g}$, and $\Phi_{1},...,\Phi_{p}$ by $\hat{\Phi}_{1},...,\hat{\Phi}_{p}$. The mode (median) of the predictive density provides the point forecasts and the quantiles of the estimated predictive density can be used to obtain estimated prediction intervals at horizon 1 (see, Online Appendix B).
4.2 Estimated Prediction Interval Uncertainty
The estimated model parameters and residuals $\hat{\varepsilon}_{t}$ can be used to build oos predictions and prediction intervals conditional on given values of the last observations in the sample, called the conditional prediction interval. In finite sample, there is uncertainty on the estimated prediction intervals resulting from semi-parametric and non-parametric estimators with different convergence rates. The estimation errors of the scalar and functional parameters have a nonlinear effect on this uncertainty, which should not be disregarded.
To highlight the specificities of the nonlinear dynamic model analysis, we compare our approach with the standard practice of predicting from a linear AR(1) model defined below:
Example 1: Linear AR(1) model: The model is given by: $y_{t}=\phi y_{t-1}+\epsilon_{t},\;|\phi|<1$ where the errors $\epsilon_{t}$ are i.i.d. with the unknown true density $f_{0}$.
Below, we compare the true and estimated prediction intervals, and introduce the bootstrap-adjusted prediction intervals.
4.2.1 True and Estimated Prediction Intervals
i) True Prediction Interval: For ease of exposition, let us consider the VAR(1) model, a forecast horizon $h=1$ and a future value of the first component series $Y_{1,T+1}$ to be forecast at date $T$ out of sample (oos) given $Y_{T}=(y_{1,T},y_{2,T})^{\prime}\equiv y$, where $(Y_{2,t})$ contains the remaining components of the series. Then, the true prediction interval at level $1-\alpha_{1}$ for $Y_{1,T+1}$ is: